# Reborn: Reinforcement-Learned Boundary Segmentation with Iterative Training for Unsupervised ASR

### v1

@@ C.1 Segmentaiton model
C.2 Hyperparameter Searches
We search the optimal hyperparameters of our phoneme predictor and segmentation models with unsupervised validation metrics. For the phoneme predictor, we directly adopt the search space of the hyperparameters and the unsupervised validation metric from wav2vec-U (Baevski et al., 2021).We search the hyperparameters indicated as $\lambda$, $\gamma$, and $\eta$ in Appendix A only during the initialization stage of the phoneme predictor (§ 3.3). We directly adopt the same hyperparameters n the following iterations of the phoneme predictor GAN training. As for the segmentation model, we take the idea from Baevski et al. (2021) and derive a similar unsupervised metric.
C.2.1   Unsupervised Validation Metric for Segmentation Model Selection
@@
Additionally, the lengths of transcriptions are taken into account. We employ NLL without length normalization, thereby favoring predictions that score high under the language model but without an undue preference for overly lengthy transcriptions.
The optimization objective is formally expressed as:
@@
captures these considerations and serves as a comprehensive measure for the performance and generalization capabilities of our unsupervised ASR model.
C.2.2   Training Details
@@
It is noteworthy that, as the iterations progress, the length constraints become less critical. This observation suggests that the model gradually learns to generate sequences of appropriate length, reducing the need for explicit length-based adjustments.
@@
Dataset LibriSpeech TIMIT MLS iter. 1 iter. 2-5 iter. 1-3 de nl fr es it pt $c_{\text{ppl}}$ 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 $c_{\text{edit}}$ 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 $c_{\text{len}}$ 0.2 0.0 0.0 0.4 0.2 0.8 0.4 0.6 0.4
C.3 WFST Decoding for Obtaining Word-Level Outputs
@@
In Reborn, phoneme-level outputs can be generated directly with greedy decoding. However, obtaining word-level outputs requires additional modules such as WFST or external language models (Baevski et al., 2021; Liu et al., 2017; Gao et al., 2023). In this work, we incorporate Reborn with WFST to generate word-level transcriptions. Our decoding procedure mainly follows Baevski et al. (2021), which builds the WFST using PyKaldi (Can et al., 2018), and additional self-loops are introduced to mimic the CTC behavior. Moreover, two hyperparameters: the acoustic scale $a$, and the additive blank weight $v$ are added during the WFST decoding process. In wav2vec-U, the search interval of $a$ is $[0,8]$ and $v$ is tuned within $[-3,8]$. As for Reborn, we find that using a much smaller search space for both of the hyperparameters is enough for obtaining reasonable outputs. More specifically, We tune the acoustic scale in $[0.4,1.2]$ with step $0.1$ and the blank weight in $[-5,2]$ with step $1$.
@@
It is worth mentioned that we utilize the original boundary generated by the segmentation model instead of the post-processed one (§ 3.1.5) for WFST decoding. We find that the increased amount of segments directly leads to longer generator outputs, which is beneficial for WFST to generate higher quality outputs. The phenomenon is also discovered by the prior works (Baevski et al., 2021; Liu et al., 2022). In wav2vec-U (Baevski et al., 2021), the segmental feature used for WFST decoding is the one before adjacent pooling; while in wav2vec-U 2.0 (Liu et al., 2022), they simply lowers the decoding stride to generate more lengthy outputs for WFST decoding.

### v2

@@ E.3 Segmentation Model Architecture
E.4 Hyperparameter Search
We search the optimal hyperparameters of our phoneme predictor and segmentation models with unsupervised validation metrics. For the phoneme predictor, we directly adopt the search space of the hyperparameters and the unsupervised validation metric from wav2vec-U [4].We search the hyperparameters indicated as $\lambda$, $\gamma$, and $\eta$ in Appendix D only during the initialization stage of the phoneme predictor (§ 3.3). We directly adopt the same hyperparameters in the following iterations of GAN training. As for the segmentation model, we take the idea from Baevski et al. [4] and derive a similar unsupervised metric.
E.5 Unsupervised Validation Metric for Segmentation Model Selection
@@
Additionally, the lengths of transcriptions are taken into account. We employ NLL without length normalization, thereby favoring predictions that score high under the language model but without an undue preference for overly lengthy transcriptions. The optimization objective is formally expressed as:
@@
E.6 Training Details
@@
It is noteworthy that, as the iterations progress, the length constraints become less critical. This observation suggests that the model gradually learns to generate sequences of appropriate length, reducing the need for explicit length-based regularizations.
@@
Dataset LibriSpeech TIMIT MLS
iter. 1 iter. 2-5 iter. 1-3 de nl fr es it pt
$c_{\text{ppl}}$ 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0 1.0
$c_{\text{edit}}$ 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2 0.2
$c_{\text{len}}$ 0.2 0.0 0.0 0.4 0.2 0.8 0.4 0.6 0.4
As for the Stage 2 GAN training, we inherit the three weighted coefficients in Eq. 9 from the previous iteration. The initial training of wav2vec-U takes 150000 steps following the original paper. After the initialization stage, we find out that optimizing with 20000 steps for each iteration is enough for converging, taking advantage of initializing the parameters from the previous stage.
All of the experiments can be done using a single NVIDIA V100 GPU. The initial stage of wav2vec-U GAN training and the Reborn stage 1 reinforcement learning takes about 12 hours of training on an NVIDIA V100 GPU for each run. As for Reborn stage 2, since the model is not randomly initialized, it only takes about 1.5 hours of training time.
E.7 WFST Decoding for Obtaining Word-Level Outputs
@@
In Reborn, phoneme-level outputs can be generated directly with greedy decoding. However, obtaining word-level outputs requires additional modules such as WFST or external language models [4, 32, 17]. In this work, we incorporate Reborn with WFST to generate word-level transcriptions. Our decoding procedure mainly follows Baevski et al. [4], which builds the WFST using PyKaldi [8], and additional self-loops are introduced to mimic the CTC behavior. Moreover, two hyperparameters: the acoustic scale $a$, and the additive blank weight $v$ are added during the WFST decoding process. In wav2vec-U, the search interval of $a$ is $[0,8]$ and $v$ is tuned within $[-3,8]$. As for Reborn, we find that using a much smaller search space for both of the hyperparameters is enough for obtaining reasonable outputs. More specifically, We tune the acoustic scale in $[0.4,1.2]$ with step $0.1$ and the blank weight in $[-5,2]$ with step $1$.
@@
It is worth mentioning that we utilize the original boundary generated by the segmentation model instead of the merged one (§ 3.1.5) for WFST decoding. We find that the increased amount of segments directly leads to longer generator outputs, which is beneficial for WFST to generate higher-quality outputs. The phenomenon is also discovered by the prior works [4, 28]. In wav2vec-U [4], the segmental feature used for WFST decoding is the one before adjacent pooling; while in wav2vec-U 2.0 [28], they simply lower the decoding stride to generate more lengthy outputs for WFST decoding.
Appendix F Speech Self-supervised Learning and Acoustic Unit Discovery
Self-supervised learning (SSL) aims to explore effective ways for representation learning without labeled data. Recent speech SSL models can be categorized into the three classes based on the pre-training objective [35], including generative [29, 27, 30], contrastive [40, 2, 3], and predictive [22, 5, 12]. The SSL representation can be utilized for many different downstream tasks, such as phoneme recognition, speaker identification, or emotion recognition [58]. Recent UASR also takes advantage from speech representations from SSL models [4, 17].
Extended from self-supervised learning, acoustic unit discovery mainly focuses on deriving phoneme or word-like units or learning speech representations that retain only linguistically relevant information in an unsupervised manner [26, 39, 16, 15, 38]. The task is highly related to unsupervised phoneme/word segmentation, and the standard evaluation protocols include the phoneme/word boundary F1, frame-level ABX score [50] on ZeroSpeech Challenge  [38], or phoneme accuracy with supervised learned linear prediction head [15]. Although both acoustic unit discovery and unsupervised ASR are learned without supervision, they differ a lot in their learning objectives. As an extended topic of representation learning, acoustic unit discovery targets on learning better-segmented representations; while unsupervised ASR directly aims at solving the distribution matching problem, and may utilize high-quality speech representations to reach the goal.
Appendix G Different Schemes for Boundary Evaluation
According to Strgar and Harwath [53], the original protocol for evaluating phoneme boundaries contains some ambiguity. The issue arises from double counting when both the ground truth boundaries and the predicted boundaries fall within the same tolerance window. To address this issue, they propose a harsh (or strict) evaluation protocol that prevents double counting. The original protocol is referred to as the lenient boundary evaluation protocol. It can be assumed that studies conducted before Strgar and Harwath [53] used the lenient evaluation method, which often overestimated the quality of the predicted boundaries. We encourage interested readers to refer to the original paper for more detailed explanations.